What is the boiling point of a solution that is produced by dissolving 20g NaOH in 2L H2O if the Kb of H2O is 0.52 oC/m?

a. 100.0oC
b. 100.3 oC
c. 100.4 oC
d. 101.0 oC
1
Expert's answer
2020-04-22T10:54:53-0400

Solution.

"\Delta T = Kb \times Cm"

"Cm = \frac{n(NaOH)}{m(H2O)}"

Cm = 0.25 mol/kg

"Cm = 0.25 \times 2 = 0.50 \ mol/kg"

"\Delta T = 0.50 \times 0.52 = 0.26"

To = 100 oC

"T(fin.) = 100 + 0.26 = 100.26 ^oC = 100.3^oC"

Answer:

b